Please help me find the product
2
(x + 7) (x - 4x + 2)

Answer by stanbon(75887)

Please help me find the product

(x + 7) (x^2 - 4x + 2)
-------
= x(x^2-4x+2) + 7(x^2-4x+2)
-------
= x^3 - 4x^2 + 2x + 7x^2 - 28x + 14
-------
= x^3 +3x^2 -26x + 14
------------------
Cheers,
Stan H.
----------------